Find three consecutive odd integers such that the square of the third integer plus the product of the other two integers is 268.
Thank you

Answer by Alan3354(69443)

Find three consecutive odd integers such that the square of the third integer plus the product of the other two integers is 268.
-------------
Use n-2, n & n+2
(n+2)^2 + n(n-2) = 268
n^2+4n+4 + n^2-2n = 268
2n^2 + 2n + 4 = 268
n^2 + n - 132 = 0
n = -12, n = 11